The Dream of the Owner The founder and the owner of UniVEND -Semih- has a dream since he has started this company, to design and produce vending machines in Turkey. Now there is an opportunity for Semih to make the dream come true. A year ago, Konya Metropolitan Municipality announced that they would be buy vending machines for the new subway systems starting from September 2011. Today is January 2012! There were no vending machines in Konya until now Part 1 According to the announcement, UnlVEND saw the ty and started to produce vending machines. Pts Answers Purchasing will be made by Municpality and will be on montly basis according to the adoption of subway users to vending machines What is the total market for vending machines? How many vending machines have been sold so far? At the end of December 2011, what percentage of the market had penetrated by UniVEND? Apply a 2nd Degree Polynomial Trendline to the monthly sales data and answer the following questions What is the expected number of sales in Sept. 2012? What is the total number of sales in 2012? Report the RA2 Fit a GAMMA Distribution Curve to the market penetration with the following parameters: a·18, b-38, cumulative-TRUE; then answer the following questions. Report the Sum of Squared Errors (SSE) of the cumulative market penetration between the actual and the GAMMA. According to the GAMMA Model, how many vending machines are ex What is the projected sales for the second half of 2013 (July December)? When will monthly sales reach the inflection point? (Report the Time Period ONLY! Using the Gamma Distribution (GAMMADIST, Cumulative TRUE) and Excel Solver, find the alpha and beta values that minimizes the SSE to be sold between March 2013 and July 2013, inclusive? d) Report beta Report SSE 13 Total Subway Users People per machine Total Market: 850,000 people everyday 250 people everyday SSE 1.8 38 3,400 machine Monthly Sales 8.00 12.00 17.00 20.00 Month September October November December January
Answer #1

1.a) total subway users: 850000

Number of users that can served by a single machine: 250

Total market for vending machine: 850000/250 = 3400

b) as per the data given below, total vending machines sold till December 2011 is summation of all the values from September to December: = 8+12+17+20= 57 (assuming these are units). Total units sold = 57

c) penetration= (units sold/total market)*100

= (57/3400)*100  

= 1.67%

3 & 4 - Sorry, I am not able to give quality answers for these as these are more related to statistics than finance, so not under my domain of expertise!.

2. a) fITTING A 2ND ORDER TRENLINE : the equation will be = y= -0.25x2 + 5.35x + 2.75

So, Sale in September would be 31

b) total sales in 2012= 341

c) R2= 0.9947